A cube and a sphere each have a surface area of 150 in, find the volume of each and determine which has the larger volume
Answer by Alan3354(69443)

A cube and a sphere each have a surface area of 150 in, find the volume of each and determine which has the larger volume
------------
Cube: SA = 6s^2 = 150
s = 5
Vol = 5^3 = 125 cubic inches
============
Sphere: SA = 4pi*r^2 = 150
r^2 = 150/(4pi)
r =~ 3.455
Vol = 172.747 cu in